# AT2018fyk: Candidate Tidal Disruption Event by a (Super)massive Black Hole Binary

### v1

@@ 3.2.2 Black-body fit to the UV/optical emission
The SED of the early power-law decay UV/optical emission in TDEs can often be well described by thermal black-body emission (van Velzen et al., 2011). We therefore measure the black-body temperature and radius, as well as the luminosity, by fitting a single temperature black-body model to the UV/optical data as a function of time. For all the UV/optical observations available, we subtract the host galaxy contribution as determined from SED fitting in Wevers et al. (2021). The host galaxy subtracted fluxes are also corrected for Galactic extinction by employing the IDL code fm_unred.pro with $E(B-V)=0.01$ (Schlegel et al., 1998), assuming the Milky Way (MW) extinction curve (Gordon et al., 2003). To carry out the black-body model fitting of the UV/optical data, only observations with positive host-subtracted fluxes in more than 2 filters are used [^5] The host contribution dominates the emission in some observations, which might result in negative host subtracted flux densities, especially in optical bands.. We use the flux density at a specific wavelength, $f_{\lambda}=4\pi R_{\rm bb}^{2}D_{L}^{-2}B_{\lambda}(T_{\rm bb},\lambda)$, to do the fitting. Here, $R_{\rm bb}$ and $T_{\rm bb}$ are the black-body radius and temperature, respectively, and $B_{\lambda}(T_{\rm bb},\lambda)$ is the Planck law as a function of wavelength. We then calculate the black-body luminosity with the Stefan-Boltzmann law, $L_{\rm bb}=4\pi\sigma_{sb}R_{\rm bb}^{2}T_{\rm bb}^{4}$, where $\sigma_{sb}$ is the Stefan-Boltzmann constant.
@@
Figure 4 shows the black-body temperature, radius and luminosity as a function of time. The relatively large uncertainties are due to the fact that the peak of the black-body spectra is not sampled by the observations. The evolution of the temperature and radius are consistent with the results of Wevers et al. (2021) for both the early and the rebrightening epochs, e.g. constant temperature and slowly decreasing radius.
@@
We fit the luminosity with the equation $\log_{10}L_{\rm bb}=-\alpha\log_{10}(t-t_{0})+C$, where $t_{0}$ is the peak day and $C$ is the luminosity at $t-t_{0}=1$. Here, we set the zero point of $t$ at MJD 58369.2 days. Giving $t_{0}=0$ d for the early epochs, the best-fit parameters are $\alpha=0.65\pm 0.09$ and $C=44.82\pm 0.20$. During the rebrightening epochs, assuming $\alpha$ is the same as the early epochs, the best-fit $t_{0}$ and $C$ are $1221.92\pm 41.19$ d and $44.14\pm 0.08$, respectively. The luminosity decays as a power-law for both the early and rebrightening epochs, but with a smaller index of 0.69 when compared with the power law decay index of $w1$. Integrating the best-fit over time, we obtain the total black-body energy released for the early and rebrightening epochs; they are $1.1\times 10^{51}$ ergs (close to the estimation of $1.4\times 10^{51}$ ergs from Wevers et al. 2019) and $1.1\times 10^{50}$ ergs, respectively.
@@
If we adopt a radiation efficiency of $\eta=0.01$ (see our analysis in the next subsection), the radiated energy amounts to $0.06M_{\odot}$ and $0.006M_{\odot}$ for the early and rebrightening epochs, respectively, indicating $M_{\star}>0.13M_{\odot}$ assuming 50% of the star’s material is unbound and 50% is accreted. This is in agreement within 1 $\sigma$ with the star’s mass $M_{\star}=0.07^{+0.12}_{-0.04}M_{\odot}$ estimated from fitting the UV light curves (see Table 5). For the early epochs, the peak black-body luminosity and temperature are $10^{44.28\pm 0.75}$ erg/s (in agreement with the estimate of $(3.0\pm 0.5)\times 10^{44}$ erg/s from Wevers et al. 2019) and $10^{4.72\pm 0.17}$ K, respectively. Given the Eddington luminosity $L_{\rm Edd}\approx 3.3\times 10^{43}$ erg/s (for the case of $M_{\bullet}=2.7\times 10^{5}M_{\odot}$), the UV black-body luminosity is about 10 times the Eddington luminosity. It takes about 200 days for the UV/optical emission to decline to the Eddington luminosity. We also plot disk bolometric luminosities derived from the slim disk modeling of the Swift X-ray spectra. We fit each of the Swift spectra with the model ${\tt TBAbs\times Thcomp\times slimd}$ separately, and then calculate the unabsorbed disk bolometric luminosity with the disk parameters from the slimd component. As shown in Figure 4, the BB luminosity of the UV/optical emission is brighter than the slim disk bolometric luminosity for the first epoch.
@@ 3.2.3 The origin of the UV/optical emission
The bolometric disk luminosity of Swift 1 can be inferred either from slim disk modeling ($1.4^{+4.0}_{-0.8}\times 10^{43}$ erg/s) or from the tdediscpec (Mummery et al., 2023b) model $(2.2^{+5.3}_{-1.7}\times 10^{43}$ erg/s). These models reach statistically consistent estimates, each of which is significantly fainter than the peak UV luminosity (see also Figure 4). If the UV/optical emission originates from a reprocessing layer, the disk would not be initially bright enough to power the layer (though at later epochs this conclusion no longer holds, see Figure 4). This indicates that the initial optical/UV emission in this event comes from shock dissipation or the outer regions of an elliptical accretion disk, and the lack of significant temperature or radius evolution suggests that this may remain the optical/UV power source at later epochs. In Section 3.3.2, we will present additional evidence to support this conclusion.
@@ 3.2.4 BH mass inferred from UV/optical emission
TDEmass posits that the dissipation of orbital energy and the production of optical/UV emission occurs due to shocks near the debris apocenter. Its radiation efficiency is a function of the semimajor axis ($a_{\rm o}$) of the debris’ orbit, $\eta_{c}\propto r_{\rm g}/a_{\rm o}$. Its temperature is determined by $T^{4}=L/(2\pi a_{\rm o}^{2}\sigma_{\rm sb})$. As $a_{\rm o}$ is a function of the specific energy of the debris, TDEmass does not aim to fit the entire UV light curve but instead focuses on the peak emission, primarily constraining two key parameters: $M_{\bullet}$ and $M_{\star}$. The estimation of $M_{\bullet}$ relies heavily on the temperature observed at peak luminosity, while $M_{\star}$ primarily depends on the peak luminosity. We use the peak BB luminosity and temperature of the early epochs, $L=10^{44.28\pm 0.75}$erg/s and $T=10^{4.72\pm 0.17}$, for the fit.
@@
MOSFiT calculates the black-body UV/optical luminosity based on the mass fallback rate, assuming a time-independent efficiency ($\eta_{\rm c}$). The mass fallback rate $\dot{M}_{\rm f}(t)$, obtained from hydrodynamical simulations of TDEs (Guillochon & Ramirez-Ruiz, 2013), depends on parameters such as $M_{\bullet}$, $M_{*}$, and $\beta$. Instead of predicting the BB radius and BB temperature from a specific theoretical model, MOSFiT fits these parameters to the observed data,assuming that the BB radius decreases as a power-law of the luminosity, $R_{\rm BB}\propto(L/L_{\rm Edd})^{l}=\dot{m}^{l}\propto t^{l^{\prime}}$. This assumption aligns with predictions from certain theoretical models; for instance, elliptical disk models predict $R_{\rm BB}\propto t^{-1.1}$ (Liu et al., 2014), while cooling envelope models predict $R_{\rm BB}\propto t^{-1}$ (Metzger, 2022), although it is in disagreement with some radiation-hydrodynamics simulations (Steinberg & Stone, 2022). An additional assumption added in MOSFiT is that the luminosity does not follow the mass fallback rate directly, but only following convolution with a “viscous delay,” adding an additional free parameter to account for potential delays in the circularization process.
@@
MOSFiT sets a soft cap to prevent the UV black-body luminosity from exceeding the Eddington luminosity. However, as we have shown, a super-Eddington UV luminosity is not impossible. As a result, we remove the soft cap from the code, allowing the luminosity to exceed the Eddington luminosity. As we have shown the decay of the $w1$ magnitude as a function of time follows the mass fallback rate, a constant radiation efficiency may be correct for early UV emission of this event. The free parameters we fit are $M_{\bullet}$, $M_{\star}$, $\beta$, viscous decay time $T_{\rm v}$, $\eta_{\rm c}$, BB radius normalization factor $n_{\rm R}$, $l$, and the first fallback time $t_{0}$ (days since first detection). We restricted our fitting to the $w2$, $m2$, and $w1$ light curves from the early epochs, and we applied priors for two specific parameters: $t_{0}$ and $\eta_{c}$, with their ranges set to [-30,0] days and [$10^{-4}$, 0.4], respectively. We use the Markov–Chain–Monte–Carlo routine in MOSFiT to perform the fit and stop the calculation until it converged, with the potential scale reduction factor (PSRF) $\leq 1.1$ (Gelman & Rubin, 1992).
@@ 3.3 X-ray versus UV/optical Timing
In this section, we investigate whether there are any time lags between the X-ray and UV/optical emissions, in order to explore possible relationships or sources of the UV/optical and hard X-ray emission.
@@ 3.3.2 Cross-Correlations between UV, Soft X-ray, and Hard X-ray Light Curves
The correlation analysis between hard and soft X-ray emissions reveals a significance above the $3\sigma$ confidence level for the L2 vs. L3 case during the initial 600 days. The calculated time lag is $-0.5^{+0.9}_{-1.6}$, consistent with a lag of 0 days. This 0-time lag is also observed in the L2 vs. L3 case during the rebrightening phase and the L2 vs. L4 case in both the early and rebrightening phases. However, it is noteworthy that the significance of the correlation diminishes in the rebrightening phase and between L2 and L4, possibly due to a poor cadence during that period and variations in the NICER observation cadence. Despite this, the persistent 0-time lag suggests a correlation between the hard and soft components. This finding aligns with our analysis in Section 3.1.2, supporting the hypothesis that the hard component originates from the corona above the disk, which generates the observed soft X-ray emission. A correlation between the hard and soft X-ray emission is also shown in some AGN system, e.g. Narrow-line Seyfert 1 galaxy 1H 0707-495 (Robertson et al., 2015), IRAS 13224-3809 (Chiang et al., 2015). We also conducted the ICCF by de-trending the X-ray light curves through subtracting the best-fit values of the power law model. The results also align with the findings here, e.g. no significant correlation between $w1$ and soft- or hard X-ray emissions, and a $3\sigma$ confidence level correlation between soft and hard X-ray emissions for case of L2 vs. L3 in the initial 600 days.
@@ 4 Discussion: A TDE in a Binary SMBH system?
In this section, we interpret our previous data analysis as evidence for the production of a TDE in a binary SMBH system, and explore its implications.
@@ 4.2 Rapid UV Light Curve Decline Suggesting Accretion Cutoff
$T_{\rm m}\approx$ $1.8\times 10^{9}~{}{\rm yr}~{}\left(\frac{T_{\rm c}}{530~{}\rm d}\right)^{8/3}\left(\frac{M_{\bullet,\rm p}}{10^{7.7}M_{\sun}}\right)^{-2/3}$ (4)
@@ 4.4 Binary TDE Event Rate
An isolated BH in a galactic nucleus will tidally disrupt stars due to diffusion in orbital angular momentum space (Frank & Rees, 1976). In this picture, stars undergo a random walk in angular momentum from uncorrelated two-body scatterings, eventually reaching pericenters less than the tidal disruption radius. The outcomes of this random walk can be quantified in terms of an angular momentum “loss cone,” which is emptied on a short dynamical timescale and refilled on a slow relaxational timescale. The TDE rate in such a system is thus equal to the loss-cone feeding rate produced by the two-body relaxation process (Lightman & Shapiro, 1977; Cohn & Kulsrud, 1978). Theoretical loss cone calculations generally find average TDE rates $\sim{\rm few}\times 10^{-5}-{\rm few}\times 10^{-4}~{}{\rm yr}^{-1}$ (Magorrian & Tremaine, 1999; Wang & Merritt, 2004; Stone & Metzger, 2016; Pfister et al., 2020; Stone et al., 2020), in some tension with (lower) recent TDE rate estimates (Yao et al., 2023).
The formation and evolution of a bound SMBHB in a galaxy’s centers would significantly change TDE event rates. Shortly after the formation of a hard SMBHB, interactions between the binary and a bound stellar cusp will enhance the TDE rate to as high as 1 yr-1, through a combination of the EKL mechanism (Kozai, 1962; Lidov, 1962; Ivanov et al., 2005b) and strong three-body scatterings (Chen et al., 2011; Wegg & Nate Bode, 2011). Mockler et al. (2023) also recently showed that around the smaller black hole, a modest binary eccentricity will allow the EKL mechanism (Ford et al., 2000; Naoz, 2016b) to produce an even larger number of TDEs from the smaller secondary BH.
@@
The OpenTDE catalog currently has 98 TDEs candidates  [^9] https://tde.space. If we conservatively take $N=1$, there would be probabilities of $\approx 100\%$ and $\approx 4\%$ to detect at least 1 TDE from the primary and the secondary, respectively, of an SMBHB out of a sample of this size. EKL mechanism can produce TDEs around the secondary that are comparable to those around the primary in the case of q=0.1 (Mockler et al., 2023). Assuming that the EKL mechanism can increase the TDE rate by an order of magnitude for other q values, the probability of detecting 1 TDE caused by the secondary out of 98 samples would then increase from $4\%$ to $31\%$. Currently, there is only one TDE, SDSS J120136.02+300305.5, explained as a SMBHB TDE originating from the primary (Liu et al., 2014). The low detected rate of binary TDEs may be due to the fact that, it is not easy to distinguish the binary TDE from the single SMBH TDE, as the perturbation from the secondary is not strong enough to produce a significant difference in the light curves. However, some factors may make the observed rated of binary TDE due to secondary higher than predicted, e.g. the direct capture of the star by the primary SMBH with $M_{\bullet}>10^{7}M_{\odot}$ due to relativistic effects, would result in a substantial reduction in the rate of TDEs by the primaries in SMBHBs (Coughlin & Nixon, 2022).

### v3

@@ 3.2.2 Black-body fit to the UV/optical emission
The SED of the early power-law decay UV/optical emission in TDEs can often be well described by thermal black-body emission (van Velzen et al., 2011). We therefore measure the black-body temperature and radius, as well as the luminosity, by fitting a single-temperature black-body model to the UV/optical data as a function of time. For all the UV/optical observations available, we subtract the host galaxy contribution as determined from SED fitting in Wevers et al. (2021). The host galaxy subtracted fluxes are also corrected for Galactic extinction by employing the IDL code fm_unred.pro with $E(B-V)=0.01$ (Schlegel et al., 1998), assuming the Milky Way (MW) extinction curve (Gordon et al., 2003). To carry out the black-body model fitting of the UV/optical data, only observations with positive host-subtracted fluxes in more than 2 filters are used [^5] The host contribution dominates the emission in some observations, which might result in negative host-subtracted flux densities, especially in optical bands.. We use the flux density at a specific wavelength, $f_{\lambda}=4\pi R_{\rm bb}^{2}D_{L}^{-2}B_{\lambda}(T_{\rm bb},\lambda)$, to do the fitting. Here, $R_{\rm bb}$ and $T_{\rm bb}$ are the black-body radius and temperature, respectively, and $B_{\lambda}(T_{\rm bb},\lambda)$ is the Planck law as a function of wavelength. We then calculate the black-body luminosity with the Stefan-Boltzmann law, $L_{\rm bb}=4\pi\sigma_{sb}R_{\rm bb}^{2}T_{\rm bb}^{4}$, where $\sigma_{sb}$ is the Stefan-Boltzmann constant.
@@
Figure 4 shows the black-body temperature, radius, and luminosity as a function of time. The relatively large uncertainties are due to the fact that the peak of the black-body spectra is not sampled by the observations. The evolution of the temperature and radius is consistent with the results of Wevers et al. (2021) for both the early and the rebrightening epochs, e.g., constant temperature and slowly decreasing radius.
@@
We fit the luminosity with the equation $\log_{10}L_{\rm bb}=-\alpha\log_{10}(t-t_{0})+C$, where $t_{0}$ is the peak day and $C$ is the luminosity at $t-t_{0}=1$. Here, we set the zero point of $t$ at MJD 58369.2 days. Giving $t_{0}=0$ d for the early epochs, the best-fit parameters are $\alpha=0.65\pm 0.09$ and $C=44.82\pm 0.20$. During the rebrightening epochs, assuming $\alpha$ is the same as the early epochs, the best-fit $t_{0}$ and $C$ are $1221.92\pm 41.19$ d and $44.14\pm 0.08$, respectively. The luminosity decays as a power-law for both the early and rebrightening epochs, but with a smaller index of 0.69 when compared with the power-law decay index of $w1$. Integrating the best-fit over time, we obtain the total black-body energy released for the early and rebrightening epochs; they are $1.1\times 10^{51}$ ergs (close to the estimation of $1.4\times 10^{51}$ ergs from Wevers et al. 2019) and $1.1\times 10^{50}$ ergs, respectively.
@@
If we adopted a radiation efficiency of $\eta=0.01$ (see our analysis in the next subsection), the radiated energy amounts to $0.06M_{\odot}$, and $0.006M_{\odot}$ for the early and rebrightening epochs, respectively, indicating $M_{\star}>0.13M_{\odot}$ assuming 50% of the star’s material is unbound and 50% is accreted. This is in agreement within 1 $\sigma$ with the star’s mass $M_{\star}=0.07^{+0.12}_{-0.04}M_{\odot}$ estimated from fitting the UV light curves (see Table 5). For the early epochs, the peak black-body luminosity and temperature are $10^{44.28\pm 0.75}$ erg/s (in agreement with the estimate of $(3.0\pm 0.5)\times 10^{44}$ erg/s from Wevers et al. 2019) and $10^{4.72\pm 0.17}$ K, respectively. Given the Eddington luminosity $L_{\rm Edd}\approx 3.3\times 10^{43}$ erg/s (for the case of $M_{\bullet}=2.7\times 10^{5}M_{\odot}$), the UV black-body luminosity is about 10 times the Eddington luminosity. It takes about 200 days for the UV/optical emission to decline to the Eddington luminosity. We also plot disk bolometric luminosities derived from the slim disk modeling of the Swift X-ray spectra. We fit each of the Swift spectra with the model ${\tt TBAbs\times Thcomp\times slimd}$ separately, and then calculate the unabsorbed disk bolometric luminosity with the disk parameters from the slimd component. As shown in Figure 4, the BB luminosity of the UV/optical emission is brighter than the slim disk bolometric luminosity for the first epoch.
@@ 3.2.3 The origin of the UV/optical emission
The bolometric disk luminosity of Swift 1 can be inferred either from slim disk modeling ($1.4^{+4.0}_{-0.8}\times 10^{43}$ erg/s) or from the tdediscpec (Mummery et al., 2023) model $(2.2^{+5.3}_{-1.7}\times 10^{43}$ erg/s). These models reach statistically consistent estimates, each of which is significantly fainter than the peak UV luminosity (see also Figure 4). If the UV/optical emission originates from a reprocessing layer, the disk would not be initially bright enough to power the layer (though at later epochs this conclusion no longer holds, see Figure 4). This indicates that the initial UV/optical emission in this event comes from shock dissipation or the outer regions of an elliptical accretion disk, and the lack of significant temperature or radius evolution suggests that this may remain the UV/optical power source at later epochs. In Section 3.3.2, we will present additional evidence to support this conclusion.
@@ 3.2.4 BH mass inferred from UV/optical emission
TDEmass posits that the dissipation of orbital energy and the production of UV/optical emission occur due to shocks near the debris apocenter. Its radiation efficiency is a function of the semimajor axis ($a_{\rm o}$) of the debris’ orbit, $\eta_{c}\propto r_{\rm g}/a_{\rm o}$. Its temperature is determined by $T^{4}=L/(2\pi a_{\rm o}^{2}\sigma_{\rm sb})$. As $a_{\rm o}$ is a function of the specific energy of the debris, TDEmass does not aim to fit the entire UV light curve but instead focuses on the peak emission, primarily constraining two key parameters: $M_{\bullet}$ and $M_{\star}$. The estimation of $M_{\bullet}$ relies heavily on the temperature observed at peak luminosity, while $M_{\star}$ primarily depends on the peak luminosity. We use the peak BB luminosity and temperature of the early epochs, $L=10^{44.28\pm 0.75}$erg/s and $T=10^{4.72\pm 0.17}$, for the fit.
@@
MOSFiT calculates the black-body UV/optical luminosity based on the mass fallback rate, assuming a time-independent efficiency ($\eta_{\rm c}$). The mass fallback rate $\dot{M}_{\rm f}(t)$, obtained from hydrodynamical simulations of TDEs (Guillochon & Ramirez-Ruiz, 2013), depends on parameters such as $M_{\bullet}$, $M_{*}$, and $\beta$. Instead of predicting the BB radius and BB temperature from a specific theoretical model, MOSFiT fits these parameters to the observed data, assuming that the BB radius decreases as a power-law of the luminosity, $R_{\rm BB}\propto(L/L_{\rm Edd})^{l}=\dot{m}^{l}\propto t^{l^{\prime}}$. This assumption aligns with predictions from certain theoretical models; for instance, elliptical disk models predict $R_{\rm BB}\propto t^{-1.1}$ (Liu et al., 2014), while cooling envelope models predict $R_{\rm BB}\propto t^{-1}$ (Metzger, 2022), although it is in disagreement with some radiation-hydrodynamics simulations (Steinberg & Stone, 2022). An additional assumption added in MOSFiT is that the luminosity does not follow the mass fallback rate directly, but only follows convolution with a “viscous delay,” adding an additional free parameter to account for potential delays in the circularization process.
@@
MOSFiT sets a soft cap to prevent the UV black-body luminosity from exceeding the Eddington luminosity. However, as we have shown, a super-Eddington UV luminosity is not impossible. As a result, we remove the soft cap from the code, allowing the luminosity to exceed the Eddington luminosity. As we have shown, the decay of the $w1$ magnitude as a function of time follows the mass fallback rate, a constant radiation efficiency may be correct for early UV emission of this event. The free parameters are $M_{\bullet}$, $M_{\star}$, $\beta$, viscous decay time $T_{\rm v}$, $\eta_{\rm c}$, BB radius normalization factor $n_{\rm R}$, $l$, and the first fallback time $t_{0}$ (days since first detection). We restricted our fitting to the $w2$, $m2$, and $w1$ light curves from the early epochs, and we applied priors for two specific parameters: $t_{0}$ and $\eta_{c}$, with their ranges set to [-30,0] days and [$10^{-4}$, 0.4], respectively. We use the Markov–Chain–Monte–Carlo routine in MOSFiT to perform the fit and stop the calculation until it converged, with the potential scale reduction factor (PSRF) $\leq 1.1$ (Gelman & Rubin, 1992).
@@ 3.3 X-ray versus UV/optical Timing
In this section, we investigate whether there are any time lags between the X-ray and UV/optical emissions, to explore possible relationships or sources of the UV/optical and hard X-ray emission.
@@ 3.3.2 Cross-Correlations between UV, Soft X-ray, and Hard X-ray Light Curves
The correlation analysis between hard and soft X-ray emissions reveals a significance above the $3\sigma$ confidence level for the L2 vs. L3 case during the initial 600 days. The calculated time lag is $-0.5^{+0.9}_{-1.6}$ days, consistent with a lag of 0 days. This 0-time lag is also observed in the L2 vs. L3 case during the rebrightening phase and the L2 vs. L4 case in both the early and rebrightening phases. However, it is noteworthy that the significance of the correlation diminishes in the rebrightening phase and between L2 and L4, possibly due to poor cadence during that period and variations in the NICER observation cadence. Despite this, the persistent 0-time lag suggests a correlation between the hard and soft components. This finding aligns with our analysis in Section 3.1.2, supporting the hypothesis that the hard component originates from the corona above the disk, which generates the observed soft X-ray emission. A correlation between the hard and soft X-ray emission is also shown in some AGN system, e.g. Narrow-line Seyfert 1 galaxy 1H 0707-495 (Robertson et al., 2015), IRAS 13224-3809 (Chiang et al., 2015). We also conducted the ICCF by de-trending the X-ray light curves through subtracting the best-fit values of the power-law model. The results also align with the findings here, e.g. no significant correlation between $w1$ and soft- or hard X-ray emissions, and a $3\sigma$ confidence level correlation between soft and hard X-ray emissions for case of L2 vs. L3 in the initial 600 days.
@@ 4 Discussion: A TDE in a Binary SMBH system?
In this section, we interpret our previous data analysis as evidence for the production of a TDE in a binary SMBH system and explore its implications.
@@ 4.2 Rapid UV Light Curve Decline Suggesting Accretion Cutoff
$\approx$ $1.8\times 10^{9}~{}{\rm yr}~{}\left(\frac{T_{\rm c}}{530~{}\rm d}\right)^{8/3}\left(\frac{M_{\bullet,\rm p}}{10^{7.7}M_{\sun}}\right)^{-2/3}$ (4)
@@ 4.4 Binary TDE Event Rate
An isolated BH in a galactic nucleus will tidally disrupt stars due to diffusion in orbital angular momentum space (Frank & Rees, 1976). In this picture, stars undergo a random walk in angular momentum from uncorrelated two-body scatterings, eventually reaching pericenters less than the tidal disruption radius. The outcomes of this random walk can be quantified in terms of an angular momentum “loss cone,” which is emptied on a short dynamical timescale and refilled on a slow relaxational timescale. The TDE rate in such a system is thus equal to the loss-cone feeding rate produced by the two-body relaxation process (Lightman & Shapiro, 1977; Cohn & Kulsrud, 1978). Theoretical loss cone calculations generally find average TDE rates $\sim{\rm few}\times 10^{-5}-{\rm few}\times 10^{-4}~{}{\rm yr}^{-1}$ (Magorrian & Tremaine, 1999; Wang & Merritt, 2004; Stone & Metzger, 2016; Pfister et al., 2020; Stone et al., 2020; Teboul et al., 2024), in some tension with (lower) recent TDE rate estimates (Yao et al., 2023).
The formation and evolution of a bound SMBHB in a galaxy’s centers would significantly change TDE event rates. Shortly after the formation of a hard SMBHB, interactions between the binary and a bound stellar cusp will enhance the TDE rate to as high as 1 yr-1, through a combination of the EKL mechanism (Kozai, 1962; Lidov, 1962; Ivanov et al., 2005) and strong three-body scatterings (Chen et al., 2011; Wegg & Nate Bode, 2011). Mockler et al. (2023) also recently showed that around the smaller black hole, a modest binary eccentricity will allow the EKL mechanism (Ford et al., 2000; Naoz, 2016) to produce an even larger number of TDEs from the smaller secondary BH.
@@
The TDE Catalog currently has 116 TDEs candidates (Goldtooth et al., 2023). If we conservatively take $N=1$, there would be probabilities of $\approx 100\%$ and $\approx 4\%$ to detect at least 1 TDE from the primary and the secondary, respectively, of an SMBHB out of a sample of this size. EKL mechanism can produce TDEs around the secondary that are comparable to those around the primary in the case of q=0.1 (Mockler et al., 2023). Assuming that the EKL mechanism can increase the TDE rate by an order of magnitude for other q values, the probability of detecting 1 TDE caused by the secondary out of 116 samples would then increase from $4\%$ to $31\%$. Currently, there are two TDEs, SDSS J120136.02+300305.5 (Liu et al., 2014) and OGLE16aaa (Shu et al., 2020), explained as SMBHB TDE originating from the primary. The low detected rate of binary TDEs may be due to the fact that, it is not easy to distinguish the binary TDE from the single SMBH TDE, as the perturbation from the secondary is not strong enough to produce a significant difference in the light curves. However, some factors may make the observed rated of binary TDE due to secondary higher than predicted, e.g. the direct capture of the star by the primary SMBH with $M_{\bullet}>10^{7}M_{\odot}$ due to relativistic effects, would result in a substantial reduction in the rate of TDEs by the primaries in SMBHBs (Coughlin & Nixon, 2022).